OpenStudy (anonymous):

x^2+3x=11 Solve using quadratic formula.

OpenStudy (anonymous):

quadratic formula can only be used if the equation is in the form \[ax^2+bx+c=0\] the quadratic formula is \[x_{1,2} = \frac{-b\pm \sqrt{b^2-4ac}}{2a}\] i suggest you memorize the quadratic formula as it will prove useful in future math courses